# A comparison of pre-existing $\Lambda$CDM predictions with the abundance of JWST galaxies at high redshift

### v1

@@ 4.2 The effect of the initial stellar mass function
In Fig. 6, we present the UVLFs from $z=7$ to $z=16$ for all three IMF variations. In all cases, the UVLFs include attenuation by dust as in the fiducial version of galform. Note that the production of metals is calculated self-consistently with the chosen IMF, and is larger for more top-heavy IMFs (see Lacey et al. 2016 for more details), and so the amount of dust also depends on the IMF. In both galform models (STD and EvolFB), the extremely top-heavy IMF case (orange curves) predicts the highest abundance of galaxies and the case of a Kennicutt (1983) IMF in both SF modes (blue curves) predicts the lowest abundance. This shows that the dual mode IMF implementation in the original galform model (where the redshift dependence arises implicitly because starbursts are more common at high redshift) is already an improvement over the assumption of a single Kennicutt (1983) IMF in all SF modes. Clearly, however, the effect is not large enough to account fully for the discrepancy between the model and the observed UVLFs.
@@ 4.3 Comparisons with other studies
Previous studies have also considered these two factors as potential solutions to the puzzle of the high-redshift JWST galaxies. For example, Ferrara et al. (2023) claimed, using an empirical model, that they were able to reproduce the UVLFs up to $z\sim 14$ if galaxies at $z\gtrsim 11$ contain negligible amounts of dust, which is roughly consistent with our findings. Iocco & Visinelli (2024), also using empirical models, suggested a negligible role of dust extinction at the highest redshifts, but stated that a modification of the star formation rate to incorporate a larger fraction of luminous objects per massive halo is also required to match the observations. However, Mauerhofer & Dayal (2023), using a simple semi-analytic model tuned to match data at $z=5-9$, claimed that dust does not have a significant impact on the visibility of early galaxies at $z\gtrsim 12$ and, in fact, reported an increase in dust content with increasing redshift at a given stellar mass. Their model underpredicts the observed UVLF at $z\sim 12-18$ by a large factor. Using a simple empirical model, Wang et al. (2023) reported that bright UVLFs at $z\sim 13$ can be fitted by a model with no dust attenuation but argued that a contribution to the luminosity from Population III stars is necessary.
Yung et al. (2024) studied the effect of the IMF on the UVLFs using the Santa Cruz semi-analytic model by simply scaling up the predicted UVLFs, which are based on a Chabrier (2003)-like IMF, by a factor of several. They claimed that a top-heavy IMF with a boost of a factor of $\sim 4$ to the UV luminosities can bring the model predictions into agreement with the observations. We note that the simple scaling that they adopted does not properly capture all the effects of the IMF, such as the change in dust attenuation with different IMFs. Thus, in our work, we choose to implement a self-consistent IMF change in galform. Trinca et al. (2024) also found that a metallicity and redshift-dependent IMF can reduce the discrepancy between models and observations. Rasmussen Cueto et al. (2023), however, argued that a top-heavy IMF alone could not solve the problem of the high abundance of bright galaxies at high redshift. They claimed that although a top-heavy IMF produces more UV photons for a given amount of stellar mass, it will also give rise to stronger SN feedback, which suppresses star formation and hence keeps the UVLFs nearly unchanged compared to the results with a standard (Salpeter, 1955) IMF.
The studies just discussed have two themes in common: dust attenuation and a top-heavy IMF. Other explanations for the existence of bright galaxies at high redshift have also been proposed. For example, Dekel et al. (2023) and Li et al. (2023) claimed that an enhanced star-formation efficiency (SFE) from feedback-free starbursts (FFS) at high redshift may enhance the abundance of massive galaxies (see also Qin et al. 2023). They do not, however, address the critical point of what the present-day luminosity function of galaxies would look like in such a scenario. Sun et al. (2023) and Shen et al. (2023) claimed that scatter in UV luminosity at a fixed halo mass alone may be enough to resolve any disagreements at high redshift (see also Yung et al. 2024), although they do not explain the physical source of such scatter. Furthermore, Mirocha & Furlanetto (2023) warned that the upscattering of low-mass halos into bins of brighter galaxies will introduce additional tensions because the resulting stellar ages, masses, and spectral slopes would be much lower than indicated by observations. To solve these extra tensions requires both star formation and dust production to be more efficient than expected at $z\gtrsim 10$, which conflicts with the expectations of previous studies (e.g. Ferrara et al. 2023; Iocco & Visinelli 2024; Wang et al. 2023), as well as this work. Note that in models like galform, scatter in the UV luminosity at a given halo mass arises automatically from the physics of galaxy formation, since halos of the same mass have different assembly histories, and leading to different formation histories for the galaxies they contain.
The studies mentioned above focused exclusively on the galaxy population at high redshift and either employed empirical models (e.g. Ferrara et al. 2023; Wang et al. 2023; Iocco & Visinelli 2024) or adopted simple scalings in semi-analytic models (e.g. Yung et al. 2024). By contrast, in this work, we employ an ab initio model of galaxy formation based on physical principles rather than empirical prescriptions, which incorporates the effects of a variable IMF, dust etc. in a self-consistent way. Most importantly, our model follows galaxy formation and evolution throughout the whole of cosmic history and accounts for a wide range of observational data at all times, not just at high redshift.
5 The descendants at $z=0$ of the high-redshift JWST galaxies
@@
In Fig. 5, we saw that the EvolFB model with a decreasing dust content towards high redshift can match the high galaxy abundance measured by JWST at $z=12-14$. Thus, in this section, we trace the bright galaxies at $z=14$ in the EvolFB model to $z=0$ and study the properties of the descendants of these high-$z$ bright galaxies. The results for the STD model are similar.
We randomly select $\sim 250$ galaxies with $M_{\rm UV}<-16$ at $z=14$. In the upper left panel of Fig. 7, we show the distributions of both the full sample (grey contours) and the selected galaxies (red plus symbols) in the ($M_{\rm UV}$, $\log\,M_{\rm halo}$) plane at $z=14$. The distribution is remarkably flat in $M_{\rm UV}$, centred around a halo mass of $M_{\rm halo}\simeq 3.5\times 10^{9}h^{-1}\,\mathrm{M_{\odot}}$ with relatively small scatter. This is very close to the lower halo mass limit for neutral hydrogen to be able to cool and make stars (Benitez-Llambay & Frenk, 2020). The spread in luminosity at given halo mass reflects the stochastic nature of the mergers that trigger the starbursts along the halo merger histories. We can see that the distribution of the selected sample captures the distribution of the full sample well.
@@
In the other three panels of Fig. 7, we show (i) the relation between stellar mass and halo mass of the descendants at $z=0$ of the selected $z=14$ bright galaxies (upper right), (ii) the relation between the halo mass of the selected $z=14$ bright galaxies and their descendants at $z=0$ (lower left), and (iii) the relation between the stellar mass of the selected $z=14$ bright galaxies and their descendants at $z=0$ (lower right). As can be seen, the $z=14$ galaxies typically reside in massive halos at $z=0$, with the halo mass of the descendants spanning a wide range from $10^{11.5}\,h^{-1}\,\mathrm{M_{\odot}}$ to $10^{15}\,h^{-1}\,\mathrm{M_{\odot}}$ with a median of $\sim 10^{13.6}\,h^{-1}\,\mathrm{M_{\odot}}$. Interestingly, the stellar mass of the descendants spans an even wider range, from dwarfs of $10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$ to very massive galaxies of $10^{11.5}\,h^{-1}\,\mathrm{M_{\odot}}$.
@@
We find that although nearly all selected galaxies were central galaxies at $z=14$, about half of them become satellites (red circles) in more massive halos at $z=0$. These satellites are captured by massive halos at different redshifts, have their gas stripped, star-formation suppressed, and stellar mass growth stopped, resulting in the wide span of stellar mass at $z=0$ seen in the figure. The descendants of high-$z$ bright galaxies that are centrals at $z=0$ populate the high-mass tail of the stellar-halo mass relation of all central galaxies at $z=0$.
@@
In Fig. 8, we present the distributions of the bulge-to-total stellar mass ratio ($B/T$) for (i) the $z=14$ selected galaxies, (ii) descendants at $z=0$ which end up as central galaxies at $z=0$, and (iii) descendants at $z=0$ which end up as satellites at $z=0$. As can be seen, the selected $z=14$ galaxies are typically extremely bulge-dominated ($B/T\sim 1$). This reflects the fact that in the model, most of these galaxies are starbursts triggered by either disk instabilities or major galaxy mergers, and these processes are assumed to create new stars in the bulge and destroy any pre-existing stellar disk. At $z=0$, however, $B/T$ becomes on average lower than at $z=14$, indicating disc formation after $z=14$. Galaxies which end up as satellites are more likely to have less bulge-dominated morphology than those that end up as centrals.
@@ 6 Conclusions and Discussion
Numerous observational papers reporting the discovery of galaxies at very high redshift by JWST have claimed, with little justification, that the abundance and/or brightness of these galaxies cannot be accommodated within the $\Lambda$CDM model (see references in the Introduction). In this work, we have compared pre-existing predictions for the UV luminosity functions of high redshift galaxies to the JWST data. These predictions, based on the Durham semi-analytic model of galaxy formation, galform, applied to merger trees from the P–Millennium N-body simulation, were presented in Cowley et al. (2018) five years prior to the launch of JWST. The predictions generally agree well with the data, particularly when the model is extended to account for the fact that the timescale for the growth of dust grains is of the same order as the age of the Universe at the highest redshifts probed by JWST, $z\gtrsim 12$. We extended our original model by considering two different ways in which the dust-to-metal ratio, DTM, depends on metallicity. Our main conclusions are:
@@
    At higher redshifts, $z\sim 12-14$, the original model underpredicts the abundance of galaxies, especially at the bright end (Fig. 4). If dust attenuation is ignored, however, the model matches the data at $z=14$ and overpredicts the counts at $z=12$. Including either of the two extensions that take into account the dust growth timescale (particularly, the D-$\tau$ model) results in a close match to the UVLFs at all redshifts.
@@
    The galaxies seen by JWST at $z=14$ have stellar masses in a narrow range ($\sim 10^{6}-10^{7.5}\,h^{-1}\,\mathrm{M_{\odot}}$) and reside in halos with a narrow range of mass ($\sim 10^{9.3}-10^{10.2}\,h^{-1}\,\mathrm{M_{\odot}}$). In contrast, their descendants at $z=0$ span a wide range of stellar ($\sim 10^{7}-10^{11.5}\,h^{-1}\,\mathrm{M_{\odot}}$) and halo ($\sim 10^{12}-10^{15}\,h^{-1}\,\mathrm{M_{\odot}}$) masses. Although the JWST galaxies are all central galaxies at $z=14$, half of them have become satellites in more massive halos by the present day (Fig. 7).
@@
    The $z=14$ bright galaxies are typically bulge-dominated ($B/T\sim 1$), while their descendants at $z=0$ have significant disc components, especially those that end up as satellites.
@@ Acknowledgements
We thank Philip Wiseman for providing a thorough explanation of the metallicity-dependent dust-to-metal ratio and Zefeng Li for helpful discussions on dust growth history. We acknowledge support from STFC Consolidated Grant ST/X001075/1. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship (grant number MR/V023381/1). CSF acknowledges support by the European Research Council (ERC) through Advanced Investigator grant, DMIDAS (GA 786910). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1, and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the UK National e-Infrastructure.

### v2

@@ 4.2 The effect of the initial stellar mass function
In Fig. 9, we present the UVLFs from $z=7$ to $z=16$ for all three IMF variations. In all cases, the UVLFs include attenuation by dust as in the fiducial version of galform. Note that the production of metals is calculated self-consistently for the chosen IMF, and is larger for more top-heavy IMFs (see Lacey et al. 2016 for more details), and so the amount of dust also depends on the IMF. In both galform models (STD and EvolFB), the extremely top-heavy IMF case (orange curves) predicts the highest abundance of galaxies and the case of a Kennicutt (1983) IMF in both SF modes (blue curves) predicts the lowest abundance. This shows that the dual mode IMF implementation in the original galform model (where the redshift dependence arises implicitly because starbursts are more common at high redshift) is already an improvement over the assumption of a single Kennicutt (1983) IMF in all SF modes. Clearly, however, the effect is not large enough to account fully for the discrepancy between the model and the observed UVLFs.
@@ 4.3 Comparisons with other studies
Previous studies have also considered these two factors as potential solutions to the puzzle of the high-redshift JWST galaxies. For example, Ferrara et al. (2023) claimed, using an empirical model, that they were able to reproduce the UVLFs up to $z\sim 14$ if galaxies at $z\gtrsim 11$ contain negligible amounts of dust, which is roughly consistent with our findings. Iocco & Visinelli (2024), also using empirical models, suggested a negligible role of dust extinction at the highest redshifts, but stated that a modification of the star formation rate to incorporate a larger fraction of luminous objects per massive halo is also required to match the observations. However, Mauerhofer & Dayal (2023), using a simple semi-analytic model tuned to match data at $z=5-9$, claimed that dust does not have a significant impact on the visibility of early galaxies at $z\gtrsim 12$ and, in fact, reported an increase in dust content with increasing redshift at a given stellar mass. Their model underpredicts the observed UVLF at $z\sim 12-18$ by a large factor. Using a simple empirical model, Wang et al. (2023) reported that bright UVLFs at $z\sim 13$ can be fitted by a model with no dust attenuation but argued that a contribution to the luminosity from Population III stars is necessary (see also Ventura et al. 2024).
Yung et al. (2024) studied the effect of the IMF on the UVLFs using the Santa Cruz semi-analytic model by simply scaling up the predicted UVLFs, which are based on a Chabrier (2003)-like IMF, by a factor of several. They claimed that a top-heavy IMF with a boost of a factor of $\sim 4$ to the UV luminosities can bring the predictions into agreement with the observations. We note that the simple scaling that they adopted does not properly capture all the effects of the IMF, such as the change in dust attenuation with different IMFs. We choose to implement a self-consistent IMF change in galform. Trinca et al. (2024) also found that a metallicity and redshift-dependent IMF can reduce the discrepancy between models and observations. Rasmussen Cueto et al. (2023), however, argued that a top-heavy IMF alone could not solve the problem of the high abundance of bright galaxies at high redshift. They claimed that although a top-heavy IMF produces more UV photons for a given amount of stellar mass, it will also give rise to stronger SN feedback, which suppresses star formation and hence keeps the UVLFs nearly unchanged compared to the results with a standard (Salpeter, 1955) IMF.
The studies above have two common themes: dust attenuation and a top-heavy IMF. Other explanations for the existence of bright galaxies at high redshift have been proposed. For example, Dekel et al. (2023) and Li et al. (2023) claimed that an enhanced star formation efficiency (SFE) from feedback-free starbursts (FFS) at high redshift may enhance the abundance of massive galaxies (see also Qin et al. 2023). They do not, however, address the critical point of what the present-day luminosity function of galaxies would look like in such a scenario. Sun et al. (2023) and Shen et al. (2023) claimed that scatter in UV luminosity at a fixed halo mass alone may be enough to resolve any disagreements at high redshift (see also Yung et al. 2024), although they do not explain the physical source of such scatter. Furthermore, Mirocha & Furlanetto (2023) warned that the upscattering of low-mass halos into bins of brighter galaxies will introduce additional tensions because the resulting stellar ages, masses, and spectral slopes would be much lower than indicated by observations. To solve these extra tensions requires both star formation and dust production to be more efficient than expected at $z\gtrsim 10$, which conflicts with the expectations of previous studies (e.g. Ferrara et al. 2023; Iocco & Visinelli 2024; Wang et al. 2023), as well as this work. Note that in models like galform, scatter in the UV luminosity at a given halo mass arises automatically from the physics of galaxy formation, since halos of the same mass have different assembly histories, leading to different formation histories for the galaxies they contain.
The above studies focused exclusively on the galaxy population at high redshift and either employed empirical models (e.g. Ferrara et al. 2023; Wang et al. 2023; Iocco & Visinelli 2024) or adopted simple scalings in semi-analytic models (e.g. Yung et al. 2024). By contrast, here, we employ an ab initio model of galaxy formation based on physical principles rather than empirical prescriptions, which incorporates the effects of a variable IMF, dust etc. in a self-consistent way. Most importantly, our model follows galaxy formation and evolution throughout the whole of cosmic history and accounts for a wide range of observational data at all times, not just at high redshift.
5 The descendants at $z=0$ of the high-redshift JWST galaxies
@@
In Fig. 8, we saw that the EvolFB model with a decreasing dust content towards high redshift approximately match the high galaxy abundance measured by JWST at $z=12-14$. In this section, we trace the bright galaxies at $z=14$ in the EvolFB model to $z=0$ and study the properties of the descendants of these high-$z$ bright galaxies. The results for the STD model are similar.
To build up the sample, we first consider all the EAGLE–DMO galaxies with $M_{\rm UV}<-16$ and $M_{\rm halo}<5\times 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$ at $z=14$. Due to the larger volume of P–Millennium, we randomly select $1/512$ (the volume ratio of the two simulations) of the galaxies with $M_{\rm UV}<-16$ and $M_{\rm halo}>5\times 10^{9}\,h^{-1}\,\mathrm{M_{\odot}}$ at $z=14$ from P–Millennium to match the sample from EAGLE–DMO. Thus, we have 161 galaxies in total, covering a wide halo mass range. In the upper left panel of Fig. 10, we show the distributions of the selected galaxies (red plus symbols) in the ($M_{\rm UV}$, $\log\,M_{\rm halo}$) plane at $z=14$. Apart from the expected increasing trend of $M_{\rm halo}$ towards the UV-brighter end, we see that galaxies with larger halo masses exhibit a greater scatter in UV magnitude.
@@
The other panels of Fig. 10 show (i) the relation between stellar mass and halo mass of the $z=0$ descendants of the selected $z=14$ bright galaxies (upper right), (ii) the relation between the halo mass of the selected $z=14$ bright galaxies and their descendants at $z=0$ (lower left), and (iii) the relation between the stellar mass of the selected $z=14$ bright galaxies and their $z=0$ descendants (lower right). The $z=14$ galaxies are seen to typically reside in massive halos at $z=0$, with the halo mass of the descendants spanning a wide range, from $10^{11}\,h^{-1}\,\mathrm{M_{\odot}}$ to $10^{15}\,h^{-1}\,\mathrm{M_{\odot}}$ with a median of $\sim 2.5\times 10^{13}\,h^{-1}\,\mathrm{M_{\odot}}$, consistent with Chen et al. (2023). Interestingly, the stellar mass of the descendants spans an even wider range, from dwarfs of $3.2\times 10^{6}\,h^{-1}\,\mathrm{M_{\odot}}$ to very massive galaxies of $3.2\times 10^{11}\,h^{-1}\,\mathrm{M_{\odot}}$.
@@
We find that although nearly all selected galaxies were centrals at $z=14$ (159 out of 161 selected galaxies), over half (101 out of 161) become satellites (red circles) in more massive halos at $z=0$. These satellites are captured by massive halos at different redshifts, have their gas stripped, star formation suppressed, and stellar mass growth stopped, resulting in the wide span of stellar mass at $z=0$ seen in the figure. The descendants of bright high-$z$ galaxies that are centrals at $z=0$ populate the high-mass tail of the stellar-halo mass relation of all central galaxies at $z=0$.
@@
We present the distributions of the bulge-to-total stellar mass ratio ($B/T$) in Fig. 11 for (i) the $z=14$ selected galaxies, (ii) descendants at $z=0$ which end up as central galaxies at $z=0$, and (iii) descendants at $z=0$ which are satellites at $z=0$. As can be seen, the selected $z=14$ galaxies are typically extremely bulge-dominated ($B/T\sim 1$). This reflects the fact that in the model, most of these galaxies are starbursts triggered by either disc instabilities or major galaxy mergers, and these processes are assumed to create new stars in the bulge and destroy any pre-existing stellar disc. At $z=0$, however, $B/T$ becomes on average lower than at $z=14$, indicating disc formation after $z=14$. Galaxies which end up as satellites are more likely to have more disc-dominated morphologies than those that become centrals.
@@ 6 Conclusions and Discussion
Numerous observational papers reporting the discovery of galaxies at very high redshift by JWST have claimed, with little justification, that the abundance and/or brightness of these galaxies cannot be accommodated within the $\Lambda$CDM model (see references in the Introduction). Here, we have compared pre-existing predictions for the UV luminosity functions of high redshift galaxies to the JWST data. These predictions, based on the Durham semi-analytic model of galaxy formation, galform, applied to merger trees from the P–Millennium N-body simulation, were presented in Cowley et al. (2018) five years prior to the launch of JWST. In this work, we compare galform predictions with JWST observations, using a combination of the large-volume P–Millennium N-body simulation and high-resolution EAGLE dark matter-only (DMO) simulation. The predictions generally agree well with the data, particularly when the model is extended to account for the fact that the timescale for the growth of dust grains is of the same order as the age of the Universe at the highest redshifts probed by JWST, $z\gtrsim 12$. We extended our original model by considering two different ways for the dust-to-metal ratio, DTM, depends on metallicity. Our main conclusions are:
@@
    At higher redshifts, $z\sim 12-14$, the original model underpredicts the abundance of galaxies, especially at the bright end (Fig. 7). If dust attenuation is ignored, however, the model matches the data at $z=14$ and overpredicts it at $z=12$. Including either of the two extensions that consider the dust growth timescale (particularly, the D-$\tau$ model) results in a close match to the UVLFs at all redshifts.
@@
    The galaxies seen by JWST at $z=14$ have stellar masses in a narrow range ($\sim 10^{6}-10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$) and reside in halos with a narrow range of mass ($\sim 10^{9}-10^{10}\,h^{-1}\,\mathrm{M_{\odot}}$). In contrast, their descendants at $z=0$ span a wide range of stellar ($\sim 3.2\times 10^{6}-3.2\times 10^{11}\,h^{-1}\,\mathrm{M_{\odot}}$) and halo ($\sim 10^{11}-3.2\times 10^{14}\,h^{-1}\,\mathrm{M_{\odot}}$) masses. Although the JWST galaxies are all central galaxies at $z=14$, over half of them have become satellites in more massive halos by the present day (Fig. 10).
@@
    The $z=14$ bright galaxies are typically bulge-dominated ($B/T\sim 1$), while their descendants at $z=0$ have significant disc components, especially those that end up as satellites (Fig. 11).
@@ Acknowledgements
We thank the anonymous referee for helpful comments that led to improvements in the paper. We thank Philip Wiseman for explaining the metallicity-dependent dust-to-metal ratio and Zefeng Li for discussions on dust growth history. We acknowledge support from STFC (ST/X001075/1). SB is supported by a UK Research and Innovation Future Leaders Fellowship (MR/V023381/1). CSF acknowledges support from the European Research Council through Advanced Investigator grant DMIDAS (GA 786910). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1, and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the UK National e-Infrastructure.